40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10397     Entry time: Thu Aug 14 23:19:49 2014
Author: rana 
Type: Summary 
Category: LSC 
Subject: ETM Violin fundamental filters moved to LSC 

 We used to do violin mode and test mass body mode notches in the SUS-LSC filter modules. Now we want them balanced in the LSC and triggered by the LSC, so they're in the filter modules which go from the the LSC output matrix to the SUS.

01.png

Today, we were getting ETM violin mode ringups while doing ALS hunt and so we moved the bandstops into the LSC. I also changed the bandstop from a wide one which missed the ETMX mode to a double bandstop which gets both the ETMX and the ETMY mode. See attached image of the Bode mag.

03.png

ELOG V3.1.3-